determine if the following set of equations has infinitely many solutions or no solutions

x-y+2z+w = 4

-2x+3y=-4

x+y+z-4w=3
Answer #1

Noco. The given System is> x-ry ry +27 + W = 4 -2x + 3y = (-4). x + y +2 -4W = 3. .Now the Coeffient Matrix and Aug men ted MNow, Let w=k Where KER. Then 2 = (2+3k) y = 2K 3 7 = I-K. Then, the Solution is (2.9. 2, w) (2+3K, 2K, 1 1-K, K) So, the syst